A doctor wants to use three lots of ichthammol containing 50%, 20% and 5% ichthammol respectively. In what proportion should they be mixed to prepare a 10% ichthammol ointment?: A doctor wants to use three lots of ichthammol containing 50%, 20% and 5% ichthammol respectively. In what proportion should they be mixed to prepare a 10% ichthammol ointment?
Answer by ankor@dixie-net.com(4538)

A doctor wants to use three lots of ichthammol containing 50%, 20% and 5% ichthammol respectively. In what proportion should they be mixed to prepare a 10% ichthammol ointment?
:
With 3 different mixtures, I don't think there is a unique solution
:
Assume a convenient resulting amt like 60
:
Let x = amt of 50% stuff, and the amt of 20% of stuff
then
(60-2x) = amt of 5% stuff
:
.50x + .20x + .05(60-2x) = .10(60)
.7x + 3 - .10x = 6
.7x - .1x = 6 - 3
.6x = 3
x = 3/.6
x = 5 units of 50% and 20% stuff
then
60 - 2(5) = 50 units of 5% stuff
:
We have 5:5:50
or
1:1:10 is the ratio of the 3 mixtures to obtain the 10% result
:
Check:
.5(1) + .2(1) + .05(10) = .1(12)
.5 + .2 + .5 = 1.2